OpenStudy (anonymous):

A large emerald with a mass of 812.04 grams was recently discovered in a mine. If the density of the emerald is 2.76grams over centimeters cubed, what is the volume? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point. Answer for Blank 1:

OpenStudy (anonymous):

@DeejNG

OpenStudy (anonymous):

i got 294.22

OpenStudy (is3535):

So, 2.76 = 812.04/x 2.76x = 812.04

OpenStudy (anonymous):

wow your good

OpenStudy (is3535):

Haha sorry, miscalculation on my part. That's correct.

OpenStudy (anonymous):

mine is correct??

OpenStudy (anonymous):

360 no scoped

OpenStudy (is3535):

No, you're fine. But, being that you're rounding to the nearest hundredth, it would be 294.22

OpenStudy (anonymous):

okay thats what i got thank you!!

OpenStudy (anonymous):

can u check one more real fast?

OpenStudy (anonymous):

Find the length of the base of a square pyramid if the volume is 128 cubic inches and has a height of 6 inches. 4 inches 8 inches 16 inches 24 inches

OpenStudy (anonymous):

i said 8in

OpenStudy (is3535):

yes you are correct we are anonymous

OpenStudy (is3535):

yes ill help

OpenStudy (anonymous):

okay thank you! So 8in is correct

OpenStudy (is3535):

1/3 * x^2 * 6 = 128 see 1/3 * 6 = 2 so, we get 2 x^2 = 128 divided by 2 on both sides, get 2x^2/2 = 128/2 x^2 = 64 x = sqrt(64) = 8 is good

OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 9 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? 96π in3 27π in3 65π in3 105π in3

OpenStudy (anonymous):

would it be 96 pi3

OpenStudy (anonymous):

do you guys know a troller around here?!?!

OpenStudy (anonymous):

96pi 3in sorry

OpenStudy (anonymous):

hes trolling right know and thats my cuzin whos trolling

OpenStudy (is3535):

8 is your answer

OpenStudy (anonymous):

yes can u help me with the other i just posted on here? also??

OpenStudy (is3535):

yea

OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 9 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? 96π in3 <---- my answer 27π in3 65π in3 105π in3

OpenStudy (is3535):

tag me ok we are anonymous